• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Deur

    Log entry time 00:19:37 on August21,2001

    Entry number 71591

    keyword=Shift summary

    data taken beam half wave plate out. target half wave plate in, target at 0 degree
    2968/21968 good e- 194k/180k
    data taken beam half wave plate out. target half wave plate in, target at 90 degree
    2969/21969 good e- 178k/166k Grand total:0.344 M
    2970/21970 good e- 192k/178k Grand total:0.714 M
    2971/21971 good e- 197k/194k Grand total:1.105 M
    2972/21972 good e- 199k/187k Grand total:1.491 M
    2972321973 good e- 189k/175k Grand total:1.855 M Note: bcm problem during this run.
    2974/21974 good e- 37k/191k Grand total: 2.083M Note: roc14 disconnected at the end of the run: bad end of run for 2974.
    2975/21975 good e- 203k/188k Grand total: 2.474M
    2976/21976 good e- 54k/44k Grand total: 2.572 M Note: Lost the right arm HV.
    2977/21977 good e- 214k/199k Grand total: 2.985M
    9:42 Target polarimetry: EPR:42.5 NMR:42.0
    10:11 Done with EPR/NMR, back to data taking.
    2978/21978 good e- 170k/158k Grand total: 3.310M
    2979/21979 good e- 149k/139k Grand total: 3.598M